OpenStudy (anonymous):

factor -x^2+2x+8 please help

OpenStudy (anonymous):

3x+10

OpenStudy (anonymous):

-x^2 is same as -x(-x) negative times a negative equals a positive

OpenStudy (jdoe0001):

\(\large -x^2+2x+8\implies \begin{array}{cccllll} -(x^2&-2x&-8)\\ &\uparrow &\uparrow \\ &2-4&2\cdot -4 \end{array} \)

OpenStudy (anonymous):

thank you so there's no other parenthesis

OpenStudy (anonymous):

?

OpenStudy (ranga):

-x^2+2x+8 = -x^2 + 4x - 2x + 8 Factor the first two terms and the last two terms and simplify.